OpenStudy (anonymous):

Simplify. (15x² - 6x + 5) - (4x² - 2x - 3) + (- 6x² + 3x - 10) = A. 17x² - x - 2 B. 13x² - 5x + 2 C. 5x² - x - 2 D. 5x² - 5x - 8

OpenStudy (anonymous):

@FR5112

OpenStudy (anonymous):

Combine Like Terms- (15x^2)+(-4x^2)+(-6x^2)= 5x^2 (-6x)+(-2x)+ (3x) = -5x (5)+(-3)+(-10) = -8 the answer is D. You understand Why?